OpenStudy (anonymous):

x+2y=5 solve for y enter your answer in this format y=( )/

OpenStudy (anonymous):

g=(2h+1)/2

OpenStudy (anonymous):

do you know how to go about getting that

OpenStudy (anonymous):

thanks for the help can you tell me were to go so i can get the answers

OpenStudy (anonymous):

5g -2h -3g = 1 2g = 1 + 2h g= (1+2h)/2

OpenStudy (anonymous):

thank you